OpenStudy (superhelp101):

If 1.5 moles of copper metal reacts with 4.0 moles of silver nitrate, how many moles of silver metal can be formed, and how many moles of the excess reactant will be left over when the reaction is complete? Unbalanced equation: Cu + AgNO3 → Cu(NO3)2 + Ag

OpenStudy (superhelp101):

@FriedRice is this like the last?

OpenStudy (anonymous):

what do you mean is this like the last?

OpenStudy (superhelp101):

like the last question that you help me with

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

you have to use molar ratio to find out

OpenStudy (superhelp101):

is this correct? balanced equation: Cu + 2AgNO3 = Cu(NO3)2 + 2Ag This tells you that 1.5 moles of Cu reacts with 2 moles AgNO3 forming 2 moles silver. There will be 4.0 - 2.0 moles of excess AgNO3 left over.

OpenStudy (anonymous):

if you want to find the excess reagent you will need to balance out the equation

OpenStudy (superhelp101):

um so is what i said wrong?

OpenStudy (superhelp101):

what 3 parts?

OpenStudy (anonymous):

ok lets start from the biggining

OpenStudy (superhelp101):

ok

OpenStudy (anonymous):

ok tell me what you did

OpenStudy (superhelp101):

this: balanced equation: Cu + 2AgNO3 = Cu(NO3)2 + 2Ag This tells you that 1.5 moles of Cu reacts with 2 moles AgNO3 forming 2 moles silver. There will be 4.0 - 2.0 moles of excess AgNO3 left over.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then how many moles of Ag can be formed

OpenStudy (anonymous):

you might need to you molar ratio to find out how many silvers were formed

OpenStudy (superhelp101):

hmm. idk how to do that

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @superhelp101 this: balanced equation: Cu + 2AgNO3 = Cu(NO3)2 + 2Ag This tells you that 1.5 moles of Cu reacts with 2 moles AgNO3 forming 2 moles silver. There will be 4.0 - 2.0 moles of excess AgNO3 left over. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

Now from the balanced reaction u can see that 1 mole of copper reacts with 2 moles of silver nitrate in order to form one mole of copper nitrate and silver metal

OpenStudy (superhelp101):

yep

OpenStudy (abhisar):

So 1.5 moles of copper metal will need 2\(\times\)1.5= ?? moles of Silver nitrate

OpenStudy (superhelp101):

3 moles of silver nitrate

OpenStudy (anonymous):

\[\frac{ 4.0molesAgNO _{3} }{ 1 } * \frac{ 169.88g AgNO _{3} }{ 1moleAgNO _{3} } *\frac{ 2molesAg }{ 2molesAgNO _{3} }\]

OpenStudy (abhisar):

Good, given is 4 moles. So how much reactant will be left. (4-3=?)

OpenStudy (superhelp101):

1. oooh

OpenStudy (abhisar):

Now 1 mole of copper was forming 2 moles of silver element, so 1.5 moles will form ?? moles of silver element.

OpenStudy (abhisar):

2\(\times\) 1.5 =?

OpenStudy (superhelp101):

1mole copper/ 2 silver element = 1.5mole/ x it is 3

OpenStudy (abhisar):

Coppper acts as a \(\bf limiting~reagent\) here. Limiting reagent, also known as the "limiting reactant", is the substance which is totally consumed when the chemical reaction is complete. The amount of product formed is limited by this reagent since the reaction cannot proceed further without it.

OpenStudy (abhisar):

So the amount of product formed depends directly on the amount of limiting reagent, in this case copper and not silver nitrate.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @superhelp101 1mole copper/ 2 silver element = 1.5mole/ x it is 3 \(\huge\checkmark\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (superhelp101):

uh uh. Thank y'all!

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

Good job @friedrice !